# Deep flows transmitted by forced surface gravity waves

### v1

@@ 4. Behavior after forcing is turned off
The above analysis corroborates the computations by Longuet-Higgins & Stewart (1964) and McIntyre (1981) and offers a path towards interpreting the results from the forcing case. As the mean flow is entirely in the carrier wave drift during forcing, and the mean flow is linear at this order, we conclude that there is an equal and opposite deep flow that exactly cancels the return flow. This scenario is shown in figure 2. The forcing is subsequently turned off and the wave packet propagates away from the generation region. This is displayed in figure 3. The packet and its return flow have no momentum. In the forcing region, we are left with the dipole that was created during the forcing event. As this can only propagate under its own image (chosen to ensure that the free surface acts as a rigid lid for this flow), this will be extremely weak and the momentum is effectively localized to this forcing region.
@@
Figure 3. Left: The second order mean stream lines during the forcing event. The momentum resides entirely in the carrier wave drift. During the forcing event, the long wave response is entirely cancelled by an equal and opposite flow induced by the forcing. Right: the forcing is now turned off and the waves propagate away from the forcing region. The wave packet again has zero net momentum, agreeing with the classical result. In the region of forcing, a dipolar flow containing the momentum imparted by the forcing event.
@@
What happens as these waves approach shore? As was shown in (64), unforced deep-water wave packets have no momentum. However, Longuet-Higgins & Stewart (1964) showed that if the carrier waves are still in deep-water, but the packet is in finite depth, the wave packet possesses momentum which we denote $I_{0}$. If a wave packets starts off so that the waves and the packet are in infinitely deep-water and then impinge upon a step such that the depth changes so that the wave packet is now in finite depth, this would seem to violate conservation of momentum as the packet has experienced a change in momentum, $I_{0}$. To balance this, somewhere in the fluid there must be momentum with value $-I_{0}$. As the carrier waves are still in deep-water, the drift (and hence momentum) associated with the carrier waves does not change over the step. Therefore, the long wave is adjusting, and weakening, when it is in finite depth versus deep-water. That excess momentum is jettisoned at the step. As it can only advect weakly under its own image charge, this flow is essentially bound to the step region.
@@ 5. Conclusion
We find that during the forcing event, the fluid has momentum which is coincident with the Stokes drift of the waves. Moreover, the additional momentum that would be associated with the deep return flow is precisely cancelled by deep interior pressure gradients sustained by the atmospheric pressure forcing at the surface. But we also find, like Longuet-Higgins & Stewart (1962), that the Stokes drift and its associated deep Eulerian return flow are carried away from the generation region with the wave packet as soon as the forcing ceases and the packet begins to freely propagate. This leads to the remarkable conclusion that ultimately, the momentum transferred from the atmosphere — which is transmitted by deep pressure forces during wave generation — resides in a deep mean flow that identically opposes the deep Eulerian return flow up to the moment that forcing stops. In non-rotating, unstratified deep water, this mean flow is non-propagating and thus remains in place as the new surface gravity wave packet propagates away.
@@
Besides the modification to the momentum budget considered here, Dysthe (1979) argued that the inclusion of the long wave mean flow modifies the modulation instability growth rates and it was subsequently shown that this gives more accurate predictions that the lower order nonlinear Schrodinger equation (Melville, 1982). However, during forcing this long wave is not presented implying a modulation to the instability growth rates.
@@
A natural question is whether stratification or rotation will significantly change the picture presented here. If the atmospheric forcing is nearly impulsive, we need consider only the initial value problem in which a surface gravity wave packet is superimposed with a Lagrangian-mean flow given by the Stokes drift, without a deep Eulerian return flow — the state of the system at the conclusion of an impulsive atmospheric forcing resonant with a surface gravity wave packet. In a stratified fluid, the resulting evolution of the freely-propagating packet was analyzed by Haney & Young (2017), who showed that the deep Eulerian return flow is slightly modified by the presence of realistic oceanic stratfication and in addition weakly radiates internal waves as the packet propagates. The mean flow left behind by the atmospheric forcing is thus accordingly weakly modified by stratification, in the opposite manner, and likely also decomposes into a weak propagating internal response together with a stronger stationary vortical response. Less clear is the fate of atmospheric momentum and evolution of a freely-propagating packet in a rotating fluid. Unlike stratification, rotation likely dramatically changes the induced flow surrounding a freely-propagating packet. For example, Ursell & Deacon (1950) (see also Pollard, 1970) showed that the shallow, Stokes-drift-associated part of the wave-induced mean flow cannot persist in a rotating fluid. As a result, the circulation around a freely-propagating packet in a rotating fluid is likely much weaker. We may then speculate that the flow left behind by the packet, which contains the momentum transferred from the atmosphere, takes the form of a packet-scale near-inertial wave with an initial state given by the wave packet Stokes drift profile.
We thank Aidan Blaser and T.S. van den Bremer for useful comments. N.P. was partially supported by NSF OCE, NASA S-MODE and NASA OVWST.

### v2

@@ 4.2. The stationary forced response
The above analysis corroborates the computations by Longuet-Higgins & Stewart (1964) and McIntyre (1981) and offers a path towards interpreting the results from the forcing case, which is shown in figure 2. As the mean flow is entirely in the carrier wave drift during forcing, and the mean flow is linear at this order, we conclude that there is an equal and opposite “wave-transmitted” flow that exactly cancels $\psi_{D}$ while the forcing is active. In other words, up to the moment that the forcing ceases there exists a mean flow
(68) $\psi_{F}|_{t=T}=-\psi_{D}|_{t=T}\,,$
that contains the momentum transmitted by the pressure forcing, through the wave field, and into the fluid below.
But this is not the whole story. The pressure condition for the forced flow changes after the pressure forcing turns off, such that $p=0$ at $b=0$. As a result, $\psi_{F}$ does not satisfy the equations of motion on its own — the total wave-transmitted flow $\psi_{WT}$ that remains at the forcing site must include a fourth component such that
(69) $\psi_{WT}=\psi_{F}+\psi_{4}\,.$
The linear equations that govern $\psi_{WT}$ are written down in (36) and (37). The latter implies that
(70) $p+gy+\partial_{b}\psi_{F}=0\,.$
As $p=0$ after the forcing is turned off, the forced response requires the fluid to change the mean water level $y(b=0)$. If we consider this change to be impulsive, we have
(71) $\frac{{\rm d}y}{{\rm d}t}=\frac{y(\mathcal{T})-y(0)}{\mathcal{T}}=\frac{y}{\mathcal{T}}=-\frac{1}{g\mathcal{T}}\partial_{b}\psi_{F}\,,$
where $\mathcal{T}$ is the time scale over which the packet has propagated away. We can explicitly compute $y$ from
(72) $\left.\partial_{b}\psi_{F}\right|_{b=0}=\gamma\int_{-\infty}^{\infty}\frac{\mathcal{A}^{2}}{(a-a^{\prime})^{2}}{\rm d}a^{\prime}=-\gamma\int_{-\infty}^{\infty}\frac{(\mathcal{A}^{2})_{a^{\prime}}}{(a-a^{\prime})}{\rm d}a^{\prime}\,,$
where
(73) ${\color[rgb]{0,0,0}\gamma=\frac{4\omega^{3}T^{2}k^{2}}{\pi g}}.$
We recognize the final integral as the Hilbert transform of $\mathcal{A}^{2}_{a^{\prime}}$, which we write as $H(\mathcal{A}^{2}_{a})$. This change in mean water level sets up a circulation that may be found by linearized equations of motion. To this end, we solve
(74) ${\color[rgb]{0,0,0}x_{\tau a}-y_{\tau b}=0;\quad x_{\tau b}+y_{\tau a}=0,}$
and define a final stream function $\psi_{4}$, so that $\dot{x}=-\psi_{4b},y_{\tau}=\psi_{4a}$ and
(75) ${\color[rgb]{0,0,0}\Delta\psi_{4}=0,}$
and
(76) ${\color[rgb]{0,0,0}\left.y_{\tau}\right|_{\tau=0,b=0}=\psi_{4a}=\frac{\gamma}{g\mathcal{T}}H(\mathcal{A}^{2}_{a}).}$
$\psi_{4a}$ and $\psi_{4b}$ are now harmonic conjugates, so we may immediately write down their solution with this boundary condition:
(77) ${\color[rgb]{0,0,0}\psi_{4a}=-\frac{\gamma}{\pi g\mathcal{T}}\int_{-\infty}^{\infty}\frac{\mathcal{A}^{2}b}{(a-a^{\prime})^{2}+b^{2}}{\rm d}a^{\prime},}$
(78) ${\color[rgb]{0,0,0}\psi_{4b}=\frac{\gamma}{\pi g\mathcal{T}}\int_{-\infty}^{\infty}\frac{\mathcal{A}^{2}(a-a^{\prime})}{(a-a^{\prime})^{2}+b^{2}}{\rm d}a^{\prime}.}$
We can integrate this to find
(79) ${\color[rgb]{0,0,0}\psi_{4}=-\frac{\gamma}{\pi g\mathcal{T}}\int_{-\infty}^{\infty}\mathcal{A}^{2}\tan^{-1}\left(\frac{b}{a-a^{\prime}}\right){\rm d}a^{\prime}+C_{0},}$
where we choose the constant $C_{0}$ so that $\psi_{4}(b=0)=0$ implying this flow has no horizontal momentum.
In figure 3, we show the streamlines when the forcing is turned off and the wave packet propagates away from the generation region. The packet and its return flow have no momentum. In the forcing region, we are left with the dipole that was created during the forcing event. As this can only propagate under its own image (chosen to ensure that the free surface acts as a rigid lid for this flow), this will be extremely weak and the momentum is effectively localized to this forcing region.
@@
Figure 3. The stream lines after the forcing is turned off. The left shows the streamlines in the forcing region, which consist of the initial long wave generated by the pressure forcing, which is equal and opposite of the long wave for an unforced packet, plus a contribution from a change in mean water level that occurs to satisfy the free surface pressure condition after the forcing is turned off. Right: the streamlines for the waves that have propagated away from the forcing region. The wave packet again has zero net momentum, agreeing with the classical result. In the region of forcing, the momentum is entirely contained in the flow that the pressure disturbance generated that was equal and opposite to the long wave response.
4.3. Waves approaching finite depth
@@
What happens as these waves approach shore? As was shown in (64), unforced deep-water wave packets have no momentum. However, Longuet-Higgins & Stewart (1964) showed that if the carrier waves are still in deep-water, but the packet is in finite depth, the wave packet possesses momentum which we denote $I_{0}$. If a wave packets starts off so that the waves and the packet are in infinitely deep-water and then impinge upon a step such that the depth changes so that the wave packet is now in finite depth, this would seem to violate conservation of momentum, as the packet has experienced a change in momentum, $I_{0}$. To balance this, somewhere in the fluid there must be momentum with value $-I_{0}$. As the carrier waves are still in deep-water, the drift (and hence momentum) associated with the carrier waves does not change over the step. Therefore, the long wave is adjusting, and weakening, when it is in finite depth versus deep-water. That excess momentum is jettisoned at the step. As waves do not have momentum in deep-water, this means that a current with momentum $-I_{0}$ must be induced. This flow can only advect weakly under its own image charge, so that it is essentially bound to the step region.
@@ 5. Conclusion
We find that during the forcing event, the fluid has momentum which is coincident with the Stokes drift of the waves. Moreover, the additional momentum that would be associated with the deep return flow is precisely canceled by deep interior pressure gradients sustained by the atmospheric pressure forcing at the surface. But we also find, like Longuet-Higgins & Stewart (1962), that the Stokes drift and its associated deep Eulerian return flow are carried away from the generation region with the wave packet as soon as the forcing ceases and the packet begins to freely propagate. This leads to the remarkable conclusion that ultimately, the momentum transferred from the atmosphere — which is transmitted by deep pressure forces during wave generation — resides in a deep mean flow that identically opposes the deep Eulerian return flow up to the moment that forcing stops. In non-rotating, unstratified deep water, this mean flow is non-propagating and thus remains in place as the new surface gravity wave packet propagates away. Note, this is invariant to the free surface being a rigid lid and instead is a consequence of the fact that in deep-water irrotational wave packets do not have momentum.
@@
Besides the modification to the momentum budget considered here, Dysthe (1979) argued that the inclusion of the long wave mean flow modifies the modulation instability growth rates and it was subsequently shown that this gives more accurate predictions that the lower order nonlinear Schrodinger equation (Melville, 1982; Pizzo & Melville, 2016, 2019). However, during forcing this long wave is not presented implying a modulation to the instability growth rates.
@@
A natural question is whether stratification or rotation will significantly change the picture presented here. If the atmospheric forcing is nearly impulsive, we need consider only the initial value problem in which a surface gravity wave packet is superimposed with a Lagrangian-mean flow given by the Stokes drift, without a deep Eulerian return flow — the state of the system at the conclusion of an impulsive atmospheric forcing resonant with a surface gravity wave packet (this situation was considered in large eddy simulations with rotation by Wagner et al. (2021)). In a stratified fluid, the resulting evolution of the freely-propagating packet was analyzed by Haney & Young (2017), who showed that the deep Eulerian return flow is slightly modified by the presence of realistic oceanic stratfication and in addition weakly radiates internal waves as the packet propagates. The mean flow left behind by the atmospheric forcing is thus accordingly weakly modified by stratification, in the opposite manner, and likely also decomposes into a weak propagating internal response together with a stronger stationary vortical response. Less clear is the fate of atmospheric momentum and evolution of a freely-propagating packet in a rotating fluid. Unlike stratification, rotation likely dramatically changes the induced flow surrounding a freely-propagating packet. For example, Ursell & Deacon (1950) (see also Pollard, 1970) showed that the shallow, Stokes-drift-associated part of the wave-induced mean flow cannot persist in a rotating fluid. As a result, the circulation around a freely-propagating packet in a rotating fluid is likely much weaker. We may then speculate that the flow left behind by the packet, which contains the momentum transferred from the atmosphere, takes the form of a packet-scale near-inertial wave with an initial state given by the wave packet Stokes drift profile.
We dedicate this paper to Sean Haney. Perhaps the only thing Sean loved more than surface waves was debating their finer details with friends, usually after surf sessions. Sean provided many useful discussions on this topic prior to his untimely passing in 2021. Be careful, Sean. We also thank an anonymous referee for useful comments that have improved the manuscript. We thank Aidan Blaser and T.S. van den Bremer for helpful comments. N.P. was partially supported by NSF OCE-2219752 and 2342714 and by NASA 80NSSC19K1037 (S-MODE) and 80NSSC23K0985 (OVWST).